OpenStudy (habbababba1):

please help im tierd of posting and getting no response will fan and medal...Eva makes a few dozen cupcakes and sells them for $8 a dozen. Which function represents the relationship between the money Eva earns, m, and the number of cupcakes she sells, c? f(c) = c × 8 f(m) = c × 8 f(c) = m × 8 f(m) = m × 8 f(m) = c × m

OpenStudy (faiqraees):

Okay she earns money for each cupcake she sells right?

OpenStudy (habbababba1):

yeah

OpenStudy (faiqraees):

which means the more cupcake she sells the more money she earns right?

OpenStudy (habbababba1):

yeah

OpenStudy (faiqraees):

so we can make an equation money = k* cupcake, right?

OpenStudy (faiqraees):

where k is just a constant

OpenStudy (habbababba1):

where do u get k?

OpenStudy (habbababba1):

i see m c and f

OpenStudy (faiqraees):

that is just a constant. Like 1,2 or 3

OpenStudy (faiqraees):

I will get to it, but just answer my question now?

OpenStudy (habbababba1):

im still confused but ok

OpenStudy (habbababba1):

the money Eva earns, m, and the number of cupcakes she sells, c?

OpenStudy (faiqraees):

And each dozen of cupcake is worth 8 dollars so m = 8c, right?

OpenStudy (habbababba1):

see we know this but how do i find out the equation

OpenStudy (faiqraees):

yes

OpenStudy (faiqraees):

Okay each dozen of cupcake is worth 8 dollars. So if she sell each dozen of cupcake she gets 8 dollars. She sells two, she gets 16 dollars. From this we can see, money = 8 * cupcake

OpenStudy (habbababba1):

f(c) = c × 8 f(m) = c × 8 f(c) = m × 8 f(m) = m × 8 f(m) = c × m

OpenStudy (habbababba1):

how do we get these equations from that tho???

OpenStudy (habbababba1):

oh i get it so would it be the second 1?

OpenStudy (faiqraees):

yes

OpenStudy (habbababba1):

ok thanks man u have no idea how annoying it gets posting all these questions and ur the first to give me a response all day